Grade 9Mechanics

A particle start from rest and moves with constant acceleration then with zero acceleration and retarded with same magnitude as accelerated .the time ratio above 1:8:1the average velocity of particle is 54m/s .find max velocity attained by the particle.

Rohitsharma
9 Years agoGrade 9

1 Answer

Shourya
9 Years ago
Avgvelocity=total distance/total timeTotal time for three respective distances be 1x+ 8x+1x=9xTotal displacement=540x (using formula) 1partWhen particle starts with rest s=1/2*ax^2 (s=ut+1/2*at^2)Eq 1Part 2Displacement=ut =ax*8x=8ax^2Eq 2Part 3S=ax*x-1/2*ax^2=ax^2-1/2ax^2Eq 3Eq1+eq2+eq39ax^2Totaldisplacement=9ax^2=540xBy solving ax=60Bycommonsence max vel. =ax=60Ans =60